3x1²+3x2²+3x3²+2x1x2+2x1x3-2x2x3

1
Expert's answer
2022-01-14T11:25:08-0500

Assume the question is as follows:

Reduce the quadratic form "3 x_{1}^{2}+3 x_{2}^{2}+3 x_{3}^{2}+2 x_{1} x_{2}+2 x_{1} x_{3}-2 x_{2} x_{3}" into 'a sum of squares' by an orthogonal transformation and give the matrix of transformation.

Solution:

On comparing the given quadratic with the general quadratic "a x^{2}+b y^{2}+c z^{2}+2 f y z +2 g z x+2 h x y" , the matrix is given by

"A=\left[\begin{array}{lll}

a & h & g \\

h & b & f \\

g & f & c

\end{array}\right]=\left[\begin{array}{rrr}

3 & 1 & 1 \\

1 & 3 & -1 \\

1 & -1 & 3

\end{array}\right]"

The desired characteristic equation becomes

"|A-\lambda I|=\left|\begin{array}{ccc}

3-\lambda & 1 & 1 \\

1 & 3-\lambda & -1 \\

1 & -1 & 3-\lambda

\end{array}\right|=0"

which is a cubic in "\lambda" and has three values viz., 1,4,4.

Hence the desired canonical form i.e., 'a sum of squares' is "x^{2}+4 y^{2}+4 z^{2}" . Solving "[A-\lambda I][X]=0" for three values of "\lambda"

For "\lambda=1" , we have "\left[\begin{array}{rrr}2 & 1 & 1 \\ 1 & 2 & -1 \\ 1 & -1 & 2\end{array}\right]" "\left[\begin{array}{l}x_{1} \\ y_{1} \\ z_{1}\end{array}\right]=0

\ or \left.\quad \begin{array}{l}2 x_{1}+y_{1}+z_{1}=0 \\ x_{1}+2 y_{1}-z_{1}=0\end{array}\right\}, i.e. \frac{x_{1}}{-1-2}=\frac{y_{1}}{1+2}=\frac{z_{1}}{4-1}=k
\\
\therefore"

"\left[\begin{array}{l}

x_{1} \\

y_{1} \\

z_{1}

\end{array}\right]=\left[\begin{array}{r}

-k \\

k \\

k

\end{array}\right]=\left[\begin{array}{r}

-1 \\

1 \\

1

\end{array}\right]"

Similarly for "\lambda=4,\left[\begin{array}{rrr}-1 & 1 & 1 \\ 1 & -1 & -1 \\ 1 & -1 & -1\end{array}\right]\left[\begin{array}{l}x \\ y \\ z\end{array}\right]=0"

We have two linearly independent vectors "X_{2}=\left[\begin{array}{l}1 \\ 1 \\ 0\end{array}\right], X_{3}=\left[\begin{array}{l}1 \\ 0 \\ 1\end{array}\right]"

As the transformation has to be an orthogonal one, therefore to obtain ' P ', first divide each elements of a corresponding eigen vector by the square root of sum of the squares of its respective elements and then express as [X Y Z]

Hence the matrix of transformation, "P=\left[\begin{array}{ccc}\frac{1}{\sqrt{3}} & \frac{1}{\sqrt{2}} & \frac{1}{\sqrt{2}} \\ \frac{1}{\sqrt{3}} & \frac{1}{\sqrt{2}} & 0 \\ \frac{1}{\sqrt{3}} & 0 & \frac{1}{\sqrt{2}}\end{array}\right]"